OpenStudy (anonymous):

George Yilling bought a share of stock at 41 3/8 on Monday. On Tuesday it cost 43 1/8 per share. Find the net change from Monday to Tuesday.

OpenStudy (kropot72):

Net change in share price is the following increase in price per share: \[43\frac{1}{8}-41\frac{3}{8}=1\frac{6}{8}=1\frac{3}{4}\]